OpenStudy (anonymous):

f(x)=5x^2-1 find the inverse and any restrictions on the domain and range, showing all steps. plz help!!!

OpenStudy (anonymous):

nick could u help plz?

OpenStudy (anonymous):

y=5x^2-1 sqrt{(y+1)/1}=x y.>-1

OpenStudy (anonymous):

@ash2326 do you know how to do this? I don't remember it...at all.

OpenStudy (ash2326):

@AbbeyD16 we have \[f(x)=5x^2-1\] we need to find inverse of f(x) that's \(f^{-1}(x)\) one important thing to keep in mind \[f(f^{-1}(x))=f^{-1} (f (x))=x\] Do you get this point ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thank you so much, Ash!

OpenStudy (ash2326):

@rebeccaskell94 we've just started now let's put \(x=f^{-1} (x)\) in f(x) so \[f(x)=5x^2-1\] \[f(f^{-1} x)=5(f^{-1} x)^2-1\] @AbbeyD16 what'd be the left side?

OpenStudy (anonymous):

Good luck, Abbey! I know Ash will do a great job with helping you! :D

OpenStudy (anonymous):

im not sure :/

OpenStudy (ash2326):

@AbbeyD16 Sorry I was disconnected, the answer is in my posts. Please read them again

OpenStudy (anonymous):

u there?

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

1. Solve x in terms of f(x) \[y=5x^2-1\] \[y+1=5x^2\] \[x^2=(y+1)/5\] \[x=\sqrt{(y+1)/5}\] Thus the inverse function \[ƒ^{−1}\] is given by the formula \[f^{-1}(y)= \sqrt{(y+1)/5}\]